A metal alloy is 25% copper. Another metal alloy is 50% copper. How much of each alloy should be mixed to create 1000 grams of an alloy that is 45% copper?
Answer by stanbon(75887)

A metal alloy is 25% copper. Another metal alloy is 50% copper. How much of each alloy should be mixed to create 1000 grams of an alloy that is 45% copper?
------------
Let amount of 25% alloy be "x" ; amount of copper in it is 0.25x grams
Amount of 50% alloy is "1000-x" ;
amount of copper is 0.5(1000-x)=500-0.5x grams
Amount of mixture is 1000 grams; amount of copper is .45(1000)=450 grams
-----------
EQUATION:
copper + copper = copper
0.25x+500-0.5x=450
-0.25x=-50
x=200 grams (amount of 25% alloy)
1000-x=800 grams (amount of 50% alloy)
--------
Cheers,
Stan H.